OpenStudy (anonymous):

Simplify: \[\sqrt[3]{4^2} \]

OpenStudy (anonymous):

ANYBODY?!

OpenStudy (anonymous):

Recall that \[4^2 = (2^2)^2 = 2^4\]\[\implies \sqrt[3]{4^2} = \sqrt[3]{2^4} = \sqrt[3]{2^3\cdot 2} = 2\sqrt[3]{2}\]

OpenStudy (anonymous):

answer is ^3sqrt(2) thank you for your explanation. what i was lacking was "factoring" the exponent

OpenStudy (anonymous):

Is anybody on here?

OpenStudy (anonymous):

yup wat happened